# Reinforcement Learning Finetunes Small Subnetworks in Large Language Models

### v1

@@ Abstract
Reinforcement learning (RL) yields substantial improvements in large language models’ (LLMs) downstream task performance and alignment with human values. Surprisingly, such large gains result from updating only a small subnetwork comprising just 5%-30% of the parameters, with the rest effectively unchanged. We refer to this phenomenon as parameter update sparsity induced by RL. It is observed across all 7 widely-used RL algorithms (e.g., PPO, GRPO, DPO) and all 10 LLMs from different families in our experiments. This sparsity is intrinsic and occurs without any explicit sparsity-promoting regularizations or architectural constraints. Finetuning the subnetwork alone recovers the test accuracy, and, remarkably, produces a model nearly identical to the one obtained via full finetuning. The subnetworks from different random seeds, training data, and even RL algorithms show substantially greater overlap than expected by chance. Our analysis suggests that this sparsity is not due to updating only a subset of layers; instead, nearly all parameter matrices receive similarly sparse updates. Moreover, the updates to almost all parameter matrices are nearly full-rank, suggesting RL updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent. We conjecture that the this update sparsity can be primarily attributed to training on data that is near the policy distribution; techniques that encourage the policy to remain close to the pretrained model, such as the KL regularization and gradient clipping, have limited impact.
@@ 1 Introduction
Reinforcement learning (RL) (Sutton et al., 1998; Ouyang et al., 2022; Ziegler et al., 2020; Ramamurthy et al., 2023; Sun et al., 2024; Zhou et al., 2025) is an important post-pretraining stage for adapting large language models (LLMs) to solving complex reasoning problems (Lightman et al., 2023; Wang et al., 2025a, c; Cui et al., 2025), alignment with human values (Ouyang et al., 2022; Bai et al., 2022; Dai et al., 2023), and adherence to safety protocols (Mu et al., 2024; Huang et al., 2024; Zhang et al., 2024; Duan et al., 2024). Since these desired behaviors often differ significantly from those of the pretrained model (Ouyang et al., 2022; Bai et al., 2022; OpenAI, 2024), it is often assumed that achieving them requires substantial changes to the model’s parameters and therefore full finetuning is widely applied during RL (Cui et al., 2025; HuggingFace, 2025; Liu et al., 2025; Pan et al., 2025; Zeng et al., 2025). While RL with full finetuning is allowed to update all parameters, does it actually do so? This paper presents surprising findings and answers this question in the negative.
@@
Finding 1 is observed in all 7 widely-used RL algorithms studied, namely PPO (Schulman et al., 2017b), GRPO (Shao et al., 2024), ORPO (Hong et al., 2024), KTO (Ethayarajh et al., 2024), DPO (Rafailov et al., 2023), SimPO (Meng et al., 2024) and PRIME (Cui et al., 2025), as well as supervised finetuning with rejection sampling (Xiong et al., 2025), and 10 models in our experiments, with the subnetworks consisting of as little as 5% of the model parameters in some cases (§3). It emerges intrinsically from RL finetuning, without any explicit sparsity-promoting regularization, architectural constraint, or use of parameter-efficient training or pruning methods. Moreover, we observe a strong consistency among the subnetworks emerged under different random seeds, training data and its order, and even different RL algorithms, suggesting that the pretrained model contains a partially transferable structure that is consistent across varied training conditions (§5).
@@
Interestingly, our experiments with PRIME suggest that approximately 20% of the parameters are consistently updated and make up the subnetwork. An additional 8% receive non-zero gradients during training that cancel out, while the remaining $\sim$70parameters remain untouched throughout the entire training process. This observation motivates us to articulate the following conjecture:
@@
More formally, let $\theta_{\text{full}}$ denote the parameters after full RL finetuning from the initial model $\theta_{\text{init}}$. Define a binary mask $m\in\{0,1\}^{|\theta_{\text{init}}|}$ where $m_{i}=1$ if $(\theta_{\text{init}}-\theta_{\text{full}})_{i}\neq 0$ and 00 otherwise. We finetune a second model from $\theta_{\text{init}}$ on the same data with the same hyperparameters and number of gradient updates, but, at each step, mask the gradients as $m\odot\nabla_{\theta}\mathcal{L}(\theta)$ right before the parameter update, so that only the identified subnetwork receives non-zero gradients. Let $\theta_{\text{sub}}$ denote the resulting parameters of this subnetwork finetuning process. Conjecture 1 states that $\theta_{\text{sub}}\approx\theta_{\text{full}}$. We provide supporting evidence for the conjecture on PRIME and DPO in §4.
@@
We find that the updates of RL finetuning do not concentrate in specific layers or components of the transformer. Instead, nearly all parameter matrices receive similarly sparse updates (§3). An exception is the layer normalization layers, which receives little to no updates. Moreover, despite the sparsity in the updates, they are almost always full-rank. This suggests that, instead of forcing the updates to reside in a low-rank subspace as in LoRA (Hu et al., 2022), RL full finetuning updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent.
@@
To better understand the potential reasons for this phenomenon, we conduct a series of experiments in §6. The results indicate that a primary factor is training on data that is near the policy distribution through, e.g., on-policy RL or performing supervised finetuning (SFT) on the same data before RL (Wang et al., 2024a; Cui et al., 2025). Intuitively, it requires less change to the policy distribution when the model learns on a sequence sampled from a distribution close to itself. In contrast, SFT often involves distribution shifts (Zhang et al., 2025) and densely updates the models in our experiments (§3). Other factors like KL-divergence regularization towards the reference model, gradient clipping (as used in PPO, GRPO, and PRIME), online vs. offline RL, all have limited impact on the sparsity of accumulated updates.
@@
Our findings have profound implications on RL finetuning of LLMs. They suggest that RL concentrates optimization entirely on a small and consistently active subnetwork, while the remaining parameters remain effectively inert. Conjecture 1 goes beyond the lottery ticket hypothesis (LTH) (Frankle and Carbin, 2019): not only can the subnetwork, finetuned in isolation, match the performance of the full model in performance as posited by LTH, we show that it also converges to an effectively identical model. These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities compared to SFT (Chu et al., 2025; Setlur et al., 2025), potentially by updating substantially fewer parameters. They also open up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (Chen et al., 2022).
@@ 2.1 Related Work
The Lottery Ticket Hypothesis (LTH; Frankle and Carbin, 2019) posited that dense neural networks contain sparse subnetworks capable of matching the performance of the full model when trained in isolation. Subsequent extensions to LLMs identified task-specific subnetworks that mitigate catastrophic forgetting without retraining entire models (Panda et al., 2024; Panigrahi et al., 2023; Yadav et al., 2023). Related efforts further discovered sparse subnetworks in pretrained language models crucial for encoding specific knowledge (Marks et al., 2025; Bayazit et al., 2024; Liu et al., 2022). Recent works have also explored exploiting the winning lotteries to improve training efficiency (Chen et al., 2022). While our observation is closely related to LTH, it differs in three core dimensions: (1) LTH identifies winning tickets by pruning, while we study subnetwork that naturally emerge; (2) LTH showed that the final model’s performance can be reproduced, we show that, in addition to the performance, the exact same model can almost be recovered; (3) LTH focuses on models trained from scratch, while we focus on finetuning from pretrained LLMs.
Sparse training methods exhibit notable benefits in RL efficiency (Graesser et al., 2022; Sidahmed et al., 2024; Sokar et al., 2022; Tan et al., 2023; Davelouis et al., 2025). Recent studies also employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) in RL and have achieved competitive performance alongside significantly reduced computational overhead (Sidahmed et al., 2024; Wang et al., 2025b). In contrast to approaches like LoRA that explicitly constrain updates to a small number of parameters, we find that fine-tuning the naturally emerging subnetwork can match or even surpass the performance of full-model finetuning. Moreover, despite its sparsity, the updates are nearly full-rank.
@@ 2.2 Background
KL-divergence regularization and gradient clipping in RL. Two widely adopted techniques to keep the policy from deviating too far from the reference model are KL-divergence regularization and gradient clipping. KL regularization (Schulman et al., 2017a), formally computed as $D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\mathbb{E}_{\pi_{\theta}}\left[\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right]$, constrains policy shifts. Gradient clipping further stabilizes training by bounding the update norm. Both are widely used in algorithms such as PPO, GRPO, and PRIME. In §6, we show that, counterintuitively, both have limited impact on the update sparsity.
@@ 3 RL Induces Sparse but Full-rank Updates; SFT Induces Dense Ones
Setup. To answer this question, we analyze publicly released model checkpoints on Hugging Face released by the authors. With the exception of models where RL is applied directly to the pretrained base model (e.g., DeepSeek-R1-Zero), most models follow a conventional three-stage pipeline: pretraining, supervised fine-tuning (SFT), and RL. We analyze both the RL and SFT stages by measuring the update sparsity between model checkpoints before and after RL or SFT fine-tuning. Our experiments cover Tulu 8B/70(Lambert et al., 2025), Eurus 7B (Yuan et al., 2025; Cui et al., 2025), DeepSeek Math 7B (Shao et al., 2024), and KTO/SimPO models (Meng et al., 2024).
@@
Results. As shown in Table 1, for all RL-finetuned models, 68.5%–96.0% of parameters remain unchanged after RL. This trend holds across different RL algorithms and model families. Particularly, Deepseek-R1-Zero presents a update sparsity of 86.0%, regardless of directly training from the pretrained base model, namely RL-Zero (DeepSeek-AI et al., 2025), and large-scale training for over 8K steps. Although exact training configurations are not always available, we observe that within the same model family, larger models tend to show higher sparsity. Importantly, all of these models are trained using full finetuning without any sparsity-promoting regularization techniques or constraints. This suggests that the update sparsity is intrinsic to RL training rather than a result of design artifacts.
In contrast, Figure 1 shows that SFT induces dense updates (only 6%-15% sparsity). These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities than SFT (Chu et al., 2025; Setlur et al., 2025), possibly by updating substantially fewer parameters.
Takeaway 1 RL leads to consistently sparse parameter updates (often $>$70% sparsity) while SFT produces dense updates. Sparsity emerges as a characteristic property of RL finetuning in LLMs.
@@
Almost all transformer layers receive similarly sparse updates. We next examine how parameter updates in RL are distributed across the model layers and individual parameter matrices (e.g., Q, K, V projections), based on DPO and PRIME models. If updates were concentrated in a subset of layers or modules, one could exploit that structure for enhancing the efficiency (Pan et al., 2024). Figure 3 shows layerwise and per-matrix sparsity across the models. The "Average Sparsity" is over each transformer layer, while others correspond to specific parameter matrices. We observe that parameter updates are distributed across different matrices rather than localized to specific ones. Except for consistently high sparsity in layer normalization layers, most layers exhibit similar sparsity levels. Our results show that sparsity is relatively even across the model. This suggests that recovering the behavior of the fully finetuned model requires updating all layers, albeit with only a subset of parameters in each.
@@ 4 Finetuning the Subnetwork Alone Can Reproduce the Full-finetuned Model
Setup. We follow the procedure described in §1 to obtain two models: one with full finetuning $\theta_{\text{full}}$, and another finetuned on the same data and hyperparameters but updating only the subnetwork $\theta_{\text{sub}}$. We experiment on two very different algorithms DPO, an off-policy algorithm using implicit outcome rewards, and PRIME, an on-policy one with process reward models, to ensure that our conclusion can generalize. We implement DPO with Open-Instruct and PRIME with verl. The exact hyperparameter choices for both can be found in Appendix B. For evaluation, we choose a subset of tasks reported in the original papers for both. For DPO we choose the LSAT (Wang et al., 2022), LogiQA (Liu et al., 2021) splits from AGIEval (Zhong et al., 2024), Math split of MMLU Pro (Wang et al., 2024b). For PRIME, we report results on the MATH500 (Hendrycks et al., 2021) benchmark across difficulty levels. For evaluation in DPO we use olmes. [^2] Open-Instruct: https://github.com/allenai/open-instruct; verl: https://github.com/volcengine/verl/; olmes: https://github.com/allenai/olmes/
@@
Results. In DPO, 94.0% weights are same between $\theta_{\text{full}}$ and $\theta_{\text{sub}}$; it is 90.5% for PRIME. Notably, for both DPO and PRIME, $\theta_{\text{full}}$ and $\theta_{\text{sub}}$ are 100% identical when using a tolerance of $10^{-4}$ instead of the default $10^{-5}$, indicating that the two models converge to nearly identical parameter values. As shown in Tables 3(a) and 3(b), $\theta_{\text{sub}}$ matches or outperforms $\theta_{\text{full}}$ on all tasks across both algorithms. These results suggest that the parameters outside the subnetwork play little role in the optimization process, and freezing them has a negligible or even beneficial impact on the model’s performance. We further observe that the training loss is consistently lower in the subnetwork finetuning setting than full finetuning (Fig. 4). They also provide supporting evidence for our Conjecture 1 in §1. This finding opens up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (Chen et al., 2022).
@@
6cm][t]0.48 Task $\theta_{\text{full}}$ $\theta_{\text{sub}}$ $\Delta$ AGIEval LSAT-AR 21.3 24.8 +3.5 AGIEval LSAT-LR 53.1 54.7 +1.6 AIGEval LogiQA-en 43.5 45.5 +2.0 GPQA 32.8 32.8 +0.0 MMLU Pro Math 50.8 51.6 +0.8 Avg 40.3 41.9 +1.6

### v2

@@ Abstract
Reinforcement learning (RL) yields substantial improvements in large language models’ (LLMs) downstream task performance and alignment with human values. Surprisingly, such large gains result from updating only a small subnetwork comprising just 5%-30% of the parameters, with the rest effectively unchanged. We refer to this phenomenon as parameter update sparsity induced by RL. It is observed across all 7 widely-used RL algorithms (e.g., PPO, GRPO, DPO) and all 10 LLMs from different families in our experiments. This sparsity occurs without any explicit sparsity-promoting regularizations or architectural constraints. Finetuning the subnetwork alone recovers the test accuracy, and, remarkably, produces a model nearly identical to the one obtained via full finetuning. The subnetworks from different random seeds, training data, and even RL algorithms show substantially greater overlap than expected by chance. Our analysis suggests that this sparsity is not due to updating only a subset of layers; instead, nearly all parameter matrices receive similarly sparse updates. Moreover, the updates to almost all parameter matrices are nearly full-rank, suggesting RL updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent. We conjecture that the this update sparsity can be primarily attributed to training on data that is near the policy distribution; techniques that encourage the policy to remain close to the pretrained model, such as the KL regularization and gradient clipping, have limited impact. Our code is available at https://github.com/SagnikMukherjee/sparsity_in_rl.
@@ 1 Introduction
Reinforcement learning (RL) (sutton1998reinforcement; ouyang2022traininglanguagemodelsfollow; ziegler2020finetuninglanguagemodelshuman; ramamurthy2023is; sun-etal-2024-aligning; zhou2025reinforcedmllmsurveyrlbased) is an important post-pretraining stage for adapting large language models (LLMs) to solving complex reasoning problems (lightman2023letsverifystepstep; wang2025otcoptimaltoolcalls; wang2025reinforcementlearningreasoninglarge; cui2025processreinforcementimplicitrewards), alignment with human values (ouyang2022traininglanguagemodelsfollow; bai2022constitutionalaiharmlessnessai; dai2023saferlhfsafereinforcement), and adherence to safety protocols (mu2024rulebasedrewardslanguage; huang2024oneshot; zhang2024negative; DBLP:journals/corr/abs-2403-03419). Since these desired behaviors often differ significantly from those of the pretrained model (ouyang2022traininglanguagemodelsfollow; bai2022constitutionalaiharmlessnessai; OpenAIOS), it is often assumed that achieving them requires substantial changes to the model’s parameters and therefore full finetuning is widely applied during RL (cui2025processreinforcementimplicitrewards; openr1; liu2025understanding; tinyzero; zeng2025simplerlzooinvestigatingtamingzero).
@@
While RL with full finetuning is allowed to update all parameters, does it actually do so? This paper presents surprising findings and answers this question in the negative. Finding 1 is observed in all 7 widely-used RL algorithms studied, namely PPO (schulman2017proximalpolicyoptimizationalgorithms), GRPO (shao2024deepseekmathpushinglimitsmathematical), ORPO (hong2024orpomonolithicpreferenceoptimization), KTO (ethayarajh2024ktomodelalignmentprospect), DPO (rafailov2023direct), SimPO (meng2024simposimplepreferenceoptimization) and PRIME (cui2025processreinforcementimplicitrewards), as well as supervised finetuning with rejection sampling (xiong2025minimalistapproachllmreasoning), and 10 models in our experiments, with the subnetworks consisting of as little as 5% of the model parameters in some cases (§3). It emerges without any explicit sparsity-promoting regularization, architectural constraint, or use of parameter-efficient training or pruning methods. Moreover, we observe a strong consistency among the subnetworks emerged under different random seeds, training data and its order, and even different RL algorithms, suggesting that the pretrained model contains a partially transferable structure that is consistent across varied training conditions (§5).
@@
Interestingly, our experiments with PRIME suggest that approximately 20% of the parameters are consistently updated and make up the subnetwork. An additional 8% receive non-zero gradients during training that cancel out, while the remaining $\sim$70% parameters remain untouched throughout the entire training process. This observation motivates us to articulate the following conjecture:
@@
More formally, let $\theta_{\text{full}}$ denote the parameters after full RL finetuning from the initial model $\theta_{\text{init}}$. Define a binary mask $m\in\{0,1\}^{|\theta_{\text{init}}|}$ where $m_{i}=1$ if $(\theta_{\text{init}}-\theta_{\text{full}})_{i}\neq 0$ and 0 otherwise. We finetune a second model from $\theta_{\text{init}}$ on the same data with the same hyperparameters and number of gradient updates, but, at each step, mask the gradients as $m\odot\nabla_{\theta}\mathcal{L}(\theta)$ right before the parameter update, so that only the identified subnetwork receives non-zero gradients. Let $\theta_{\text{sub}}$ denote the resulting parameters of this subnetwork finetuning process. Conjecture 1 states that $\theta_{\text{sub}}\approx\theta_{\text{full}}$. We provide supporting evidence for the conjecture on PRIME and DPO in §4.
@@
We find that the updates of RL finetuning do not concentrate in specific layers or components of the transformer. Instead, nearly all parameter matrices receive similarly sparse updates (§3). An exception is the layer normalization layers, which receives little to no updates. Moreover, despite the sparsity in the updates, they are almost always full-rank. This suggests that, instead of forcing the updates to reside in a low-rank subspace as in LoRA (hu2022lora), RL full finetuning updates a small subset of parameters that nevertheless span almost the full subspaces that the parameter matrices can represent.
@@
To better understand the potential reasons for this phenomenon, we conduct a series of experiments in §6. The results indicate that a primary factor is training on data that is near the policy distribution through, e.g., on-policy RL or performing supervised finetuning (SFT) on the same data before RL (wang2024mathshepherdverifyreinforcellms; cui2025processreinforcementimplicitrewards). Intuitively, it requires less change to the policy distribution when the model learns on a sequence sampled from a distribution close to itself. In contrast, SFT often involves distribution shifts (zhang2025bestinstructiontuningdatafit) and densely updates the models in our experiments (§3). Other factors like KL-divergence regularization towards the reference model, gradient clipping (as used in PPO, GRPO, and PRIME), online vs. offline RL, all have limited impact on the sparsity of accumulated updates.
@@
Our findings have important implications for the RL fine-tuning stage of LLMs. They suggest that when RL fine-tuning is performed on data closely aligned with the current policy, as is typical in practice, optimization concentrates primarily on a small, consistently active subnetwork, leaving most other parameters effectively inert. Conjecture 1 goes beyond the lottery ticket hypothesis (LTH) (frankle2018the): not only can the subnetwork, finetuned in isolation, match the performance of the full model in performance as posited by LTH, we show that it also converges to an effectively identical model. These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities compared to SFT (chu2025sftmemorizesrlgeneralizes; setlur2025scalingtesttimecomputeverification), potentially by updating substantially fewer parameters. They also open up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (chen2022coarseninggranularitystructurallysparse).
@@ 2.1 Related Work
The Lottery Ticket Hypothesis (LTH; frankle2018the) posited that dense neural networks contain sparse subnetworks capable of matching the performance of the full model when trained in isolation. Subsequent extensions to LLMs identified task-specific subnetworks that mitigate catastrophic forgetting without retraining entire models (panda2024lotteryticketadaptationmitigating; pmlr-v202-panigrahi23a; yadav2023tiesmergingresolvinginterferencemerging). Related efforts further discovered sparse subnetworks in pretrained language models crucial for encoding specific knowledge (marks2025sparsefeaturecircuitsdiscovering; bayazit2024discoveringknowledgecriticalsubnetworkspretrained; liu2022winwindealsparserobust). Recent works have also explored exploiting the winning lotteries to improve training efficiency (chen2022coarseninggranularitystructurallysparse). While our observation is closely related to LTH, it differs in three core dimensions: (1) LTH identifies winning tickets by pruning, while we study subnetwork that naturally emerge; (2) LTH showed that the final model’s performance can be reproduced, we show that, in addition to the performance, the exact same model can almost be recovered; (3) LTH focuses on models trained from scratch, while we focus on finetuning from pretrained LLMs.
Sparse training methods exhibit notable benefits in RL efficiency (graesser2022statesparsetrainingdeep; sidahmed2024parameterefficientreinforcementlearning; sokar2022dynamicsparsetrainingdeep; tan2023rlx2trainingsparsedeep; davelouis2025interplaysparsitytrainingdeep). Recent studies also employ Low-Rank Adaptation (LoRA) (hu2022lora) in RL and have achieved competitive performance alongside significantly reduced computational overhead (sidahmed2024parameterefficientreinforcementlearning; wang2025tinatinyreasoningmodels). In contrast to approaches like LoRA that explicitly constrain updates to a small number of parameters, we find that fine-tuning the naturally emerging subnetwork can match or even surpass the performance of full-model finetuning. Moreover, despite its sparsity, the updates are nearly full-rank.
@@ 2.2 Background
KL-divergence regularization and gradient clipping in RL. Two widely adopted techniques to keep the policy from deviating too far from the reference model are KL-divergence regularization and gradient clipping. KL regularization (schulman2017trustregionpolicyoptimization), formally computed as $D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})=\mathbb{E}_{\pi_{\theta}}\left[\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right]$, constrains policy shifts. Gradient clipping further stabilizes training by bounding the update norm. Both are widely used in algorithms such as PPO, GRPO, and PRIME. In §6, we show that, counterintuitively, both have limited impact on the update sparsity.
@@ 3 RL Induces Sparse but Full-rank Updates; SFT Induces Dense Ones
Setup. To answer this question, we analyze publicly released model checkpoints on Hugging Face released by the authors. With the exception of models where RL is applied directly to the pretrained base model (e.g., DeepSeek-R1-Zero), most models follow a conventional three-stage pipeline: pretraining, supervised fine-tuning (SFT), and RL. We analyze both the RL and SFT stages by measuring the update sparsity between model checkpoints before and after RL or SFT fine-tuning. Our experiments cover Tulu 8B/70(lambert2025tulu3pushingfrontiers), Eurus 7B (yuan2024advancing; cui2025processreinforcementimplicitrewards), DeepSeek Math 7B (shao2024deepseekmathpushinglimitsmathematical), and KTO/SimPO models (meng2024simposimplepreferenceoptimization).
@@
Results. As shown in Table 1, for all RL-finetuned models, 68.5%–96.0% of parameters remain unchanged after RL. This trend holds across different RL algorithms and model families. Particularly, Deepseek-R1-Zero presents a update sparsity of 86.0%, regardless of directly training from the pretrained base model, namely RL-Zero (deepseekai2025deepseekr1incentivizingreasoningcapability), and large-scale training for over 8K steps. Although exact training configurations are not always available, we observe that within the same model family, larger models tend to show higher sparsity. Importantly, all of these models are trained using full finetuning without any sparsity-promoting regularization techniques or constraints. This suggests that the update sparsity emerges naturally.
In contrast, Figure 1 shows that SFT induces dense updates (only 6%-15% sparsity). These results offer fresh evidence supporting recent findings that RL better preserves the pretrained capabilities than SFT (chu2025sftmemorizesrlgeneralizes; setlur2025scalingtesttimecomputeverification), possibly by updating substantially fewer parameters.
Takeaway 1 RL leads to consistently sparse parameter updates (often >>70% sparsity) while SFT produces dense updates. This sparsity emerges without regularizations or architectural constraints.
@@
Almost all transformer layers receive similarly sparse updates. We next examine how parameter updates in RL are distributed across the model layers and individual parameter matrices (e.g., Q, K, V projections), based on DPO and PRIME models. If updates were concentrated in a subset of layers or modules, one could exploit that structure for enhancing the efficiency (pan2024lisa). Figure 3 shows layerwise and per-matrix sparsity across the models. The "Average Sparsity" is over each transformer layer, while others correspond to specific parameter matrices. We observe that parameter updates are distributed across different matrices rather than localized to specific ones. Except for consistently high sparsity in layer normalization layers, most layers exhibit similar sparsity levels. Our results show that sparsity is relatively even across the model. This suggests that recovering the behavior of the fully finetuned model requires updating all layers, albeit with only a subset of parameters in each.
@@ 4 Finetuning the Subnetwork Alone Can Reproduce the Full-finetuned Model
Setup. We follow the procedure described in §1 to obtain two models: one with full finetuning $\theta_{\text{full}}$, and another finetuned on the same data and hyperparameters but updating only the subnetwork $\theta_{\text{sub}}$. We experiment on two very different algorithms DPO, an off-policy algorithm using implicit outcome rewards, and PRIME, an on-policy one with process reward models, to ensure that our conclusion can generalize. We implement DPO with Open-Instruct and PRIME with verl. The exact hyperparameter choices for both can be found in Appendix B. For evaluation, we choose a subset of tasks reported in the original papers for both. For DPO we choose the LSAT (10.1109/TASLP.2022.3164218), LogiQA (10.5555/3491440.3491941) splits from AGIEval (zhong-etal-2024-agieval), Math split of MMLU Pro (wang2024mmlupro). For PRIME, we report results on the MATH500 (hendrycks2021measuring) benchmark across difficulty levels. For evaluation in DPO we use olmes. [^2] Open-Instruct: https://github.com/allenai/open-instruct; verl: https://github.com/volcengine/verl/; olmes: https://github.com/allenai/olmes/
@@
Results. In DPO, 94.0% weights are same between $\theta_{\text{full}}$ and $\theta_{\text{sub}}$; it is 90.5% for PRIME. Notably, for both DPO and PRIME, $\theta_{\text{full}}$ and $\theta_{\text{sub}}$ are 100% identical when using a tolerance of $10^{-4}$ instead of the default $10^{-5}$, indicating that the two models converge to nearly identical parameter values. As shown in Tables 3(a) and 3(b), $\theta_{\text{sub}}$ matches or outperforms $\theta_{\text{full}}$ on all tasks across both algorithms. These results suggest that the parameters outside the subnetwork play little role in the optimization process, and freezing them has a negligible or even beneficial impact on the model’s performance. We further observe that the training loss is consistently lower in the subnetwork finetuning setting than full finetuning (Fig. 4). They also provide supporting evidence for our Conjecture 1 in §1. This finding opens up new possibilities for more efficient RL training methods that explicitly leverage this update sparsity (chen2022coarseninggranularitystructurallysparse).
@@
6cm][t]0.48 Task $\theta_{\text{full}}$ $\theta_{\text{sub}}$ $\Delta$ AGIEval LSAT-AR 21.3 24.8 +3.5 AGIEval LSAT-LR 53.1 54.7 +1.6 AGIEval LogiQA-en 43.5 45.5 +2.0 GPQA 32.8 32.8 +0.0 MMLU Pro Math 50.8 51.6 +0.8 Avg 40.3 41.9 +1.6